OpenStudy (anonymous):

7m^2+2m=8 use quad formula, help? please? im timed as usual :(

OpenStudy (anonymous):

\[7m^2 +2m - 8 = 0\] a = 7 b = 2 c = -8\[\frac{-2 \pm \sqrt{2^2 - 4(7)(-8)}}{14}\] \[\frac{-2 \pm \sqrt{228}}{14}\]\[\frac{-2 \pm 2 \sqrt{57}}{14}\] \[\frac{-1 \pm \sqrt{57}}{7}\]

OpenStudy (anonymous):

THANKS :D :D you rock!